[Brahms-l] Shift report 20040208 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Sun Feb 08 2004 - 00:01:28 EST
------------------------------------------------------------
Date: Sun Feb  8 00:01:28 2004, Shift: 16:00-24:00
Supervisor: JHL, with: 
Log book#15 pages 50 - 50

Shift Summary:
------------------------------------------------------------
1. Beam for ~4.5 hours 
   18:00-22:00 (zdc/sec: 2100->600) 
   23:20-      (zdc/sec: 3500->   ) 

2. Data  
   FS 2.3deg 0.71B:runs 10199 (340K trig2), 10200(150k trig2)  
   FS 2.3deg 0.5B: runs 10201 (82k trig2),10202(160k+running) 
   MRS 45deg 2255A runs 10199 - 10202 (540k+running) 

3. No new detector problems 
  (BB Big tubes are still off: see e-log Feb. 6 8:00-16:00) 


------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Feb 8 00:08:52 2004

This archive was generated by hypermail 2.1.8 : Sun Feb 08 2004 - 00:09:15 EST